the sum joshua and alyssa's age is 36. josh is 4 years younger then alyssa, how old are they

let x rep. josh
let y rep. alyssa

Answer by Boreal(15235)

x+y=36
y=x+4
substitute for y
x+x+4=36
2x+4=36, collected terms, now subtract 4 from each side.
2x=32 , now divide both sides by 2.
x=16
x+4 =20
Josh is 16
Alyssa is 20.